### v1

# MASQ: Multi-Agent Reinforcement Learning for Single Quadruped Robot Locomotion

Abstract
This paper proposes a novel method to improve locomotion learning for a single quadruped robot using multi-agent deep reinforcement learning (MARL). Many existing methods use single-agent reinforcement learning for an individual robot or MARL for the cooperative task in multi-robot systems. Unlike existing methods, this paper proposes using MARL for the locomotion learning of a single quadruped robot. We develop a learning structure called Multi-Agent Reinforcement Learning for Single Quadruped Robot Locomotion (MASQ), considering each leg as an agent to explore the action space of the quadruped robot, sharing a global critic, and learning collaboratively. Experimental results indicate that MASQ not only speeds up learning convergence but also enhances robustness in real-world settings, suggesting that applying MASQ to single robots such as quadrupeds could surpass traditional single-robot reinforcement learning approaches. Our study provides insightful guidance on integrating MARL with single-robot locomotion learning.
Index Terms:
Robot control, quadruped robot locomotion, reinforcement learning (RL), multi-agent RL.
I Introduction
Deep reinforcement learning (RL) has made remarkable progress in various robot control learning, such as quadruped robot [1], biped robot [2], and unmanned aerial vehicle [3]. This paper focuses on the quadruped robot control learning.
In the process of applying deep RL control to a single robot, it is prevalent to use single-agent algorithms[4, 5, 6, 7]. However, single-agent algorithms have limits in managing coordination in specific problems. Many existing methods use single-agent RL algorithms for individual robot learning or multi-agent deep reinforcement learning (MARL) for multi-robot systems in cooperative tasks [8, 9, 10]. MARL Cooperation has been widely demonstrated in Game-AI [11, 12] to have significant advantages in multi-agent collaboration.
Figure 1: Quadruped robot on various terrains
Unlike existing approaches, we propose using MARL for the locomotion learning of a single quadruped robot to enhance cooperation between its four legs, enabling it to navigate complex terrains and perform intricate tasks. By introducing cooperation MARL, where each leg acts as an agent, the quadruped robot can better coordinate its movements. This collaborative learning structure, termed Multi-Agent Reinforcement Learning for Single Quadruped Robot Locomotion (MASQ), allows the robot to share experiences. Fig. 1 presents the deployment of the MASQ algorithm on a quadruped robot tested across different terrains, including grass, rocks, flat surfaces, and rubber tracks, demonstrating its performance in diverse environments. Fig.2 shows a sim-to-real comparison of trot gaits in a quadruped robot, highlighting the consistency between simulated and real-world gait patterns.
The main contributions of this paper can be summarized as follows:
  • This paper proposes MASQ, a method treats each leg of the single quadruped robot as an agent. The locomotion learning can be modeled and solved as a cooperative MARL problem.
  • Experimental results show that the proposed method enhances performance in executing gaits, outperforms traditional single-agent approaches and improves the efficiency in training and robustness of robotic systems.
Figure 2: Sim-to-Real Comparison of Trot Gaits
II Related Work
II-A Deep RL for single robot control
Recent advances in deep RL for quadruped robots are driven by simulation technologies such as Isaac Gym [13, 14]. Hardware advancements have shown robust performance on various tasks through Sim2Real transfer with zero shot [15, 16, 17, 18]. Current research focuses on task-specific reward composition and training paradigms to bridge the Sim2Real gap [19, 20, 21], often using Proximal Policy Optimization (PPO) [22] with an emphasis on task and reward design rather than novel RL algorithms. Although some studies have explored learning algorithms to improve efficiency [23, 24, 25, 26, 27, 28], most efforts remain centered on efficiency rather than modifying algorithms to control the characteristics of the object. Additionally, research has delved into task-specific strategies like rapid motor adaptation (RMA) [19], hierarchical RL for multi-skill tasks [29], and symmetry-based data enhancement [30, 31], yet challenges remain in using robot symmetry through single-agent methods, indicating that algorithmic research on this aspect is still underdeveloped.
II-B MARL for multi-robot control
In multi-agent settings, algorithms like Multi-Agent Proximal Policy Optimization (MAPPO) [32]. Temporally Extended Multi-Agent Reinforcement Learning (TEMP) [33] have demonstrated strong capabilities in addressing multi-agent robotic challenges, such as drone fleet control [34, 35] and autonomous vehicle fleets [36, 37]. Methods like Reinforced Inter-Agent Learning (RIAL) and Differentiable Inter-Agent Learning (DIAL) [38] further enhance collaborative performance by developing communication protocols among agents. This paper proposes modeling single-quadruped robot locomotion learning as a cooperative MARL problem, where each leg is treated as an independent agent, contrasting with previous approaches that treat the robot as a monolithic entity [39, 40, 41, 42] or cooperative groups of robots [43, 44, 45].
III Preliminaries
This paper considers a finite-horizon Markov decision process (MDP) [46], defined by a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma,T)$. $\mathcal{S}$ denotes the state space, $\mathcal{A}:=\{a_{0},a_{1},...,a_{|\mathcal{A}|-1}\}$ represents a finite action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ represents the state transition distribution, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ denotes a reward function, $\gamma\in[0,1)$ denotes a discount factor, and $T$ is a time horizon. At each time step $t$, the policy $\pi$ selects an action $a_{t}\in\mathcal{A}$. After entering the next state by sampling from $\mathcal{P}\left(s_{t+1}\mid s_{t},a_{t}\right)$, the agent receives an immediate reward $r\left(s_{t},a_{t}\right)$. The agent continues to perform actions until it enters a terminal state or $t$ reaches the time horizon $T$. Reinforcement learning (RL) aims to learn a policy $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ for decision-making problems by maximizing discounted rewards. For any policy $\pi$, the state-action value function ($Q$ function) is defined as
$Q^{\pi}(s,a)={\mathbb{E}^{\pi}}\left[\sum_{t=0}^{T}\gamma^{t}r\left(s_{t},a_{t}\right)\mid s_{0}=s,a_{0}=a\right]$ (1)
Proximal Policy Optimization (PPO) enhances the stability and performance of policy gradient methods by limiting policy updates to prevent destabilizing deviations. The core PPO update rule optimizes a clipped surrogate function:
$L^{\text{CLIP}}(\theta)=\mathbb{E}_{t}\left[\min\left(r_{t}(\theta)\hat{A}_{t},\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}\right)\right]$ (2)
where
$r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}$ (3)
and
$\hat{A}_{t}=R_{t}-V_{\phi}(s_{t})$ (4)
Policy parameters $\theta$ are updated as:
$\theta\leftarrow\theta+\alpha\nabla_{\theta}L^{\text{CLIP}}(\theta)$ (5)
PPO’s constrained updates stabilize training and improve performance, making it practical for complex single-agent RL tasks.
IV Multi-Agent Reinforcement Learning for Single Quadruped Robot Locomotion Learning
Figure 3: The framework of MASQ
Although many approaches employ deep RL to address single-agent motion control, relying solely on single-agent methods can constrain the complexity and adaptability of learned behaviors in dynamic environments. In this work, we utilize the collaborative potential of multiple agents to improve the learning process for a single robot’s locomotion, enhancing its stability, agility, and adaptability. The proposed method allows multiple virtual agents to independently explore the action space of the quadruped robot, sharing their experiences and learning collaboratively. It enables the robot to benefit from diverse strategies and insights that are difficult to achieve through a single-agent perspective. Specifically, each leg of the quadruped robot is treated as a separate agent within the multi-agent structure, with individual observations and a shared knowledge base, significantly enhancing the cooperation among the robot’s limbs for more effective locomotion.
IV-A MASQ modeling
This paper formulates quadruped robot locomotion learning as a cooperative multi-agent problem, which is described as a partially observable decentralized Markov decision process (decPOMDP) [47]. The decPOMDP is defined by the tuple $G=(\mathcal{S},\mathcal{A},\mathcal{P},r,\mathcal{Z},\mathcal{O},N,\gamma,T)$. Here, $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{P}$ is the state transition function, $r$ is the reward function, $\mathcal{Z}$ is the observation space, $\mathcal{O}$ is the observation function, $N$ is the number of agents, $\gamma$ is the discount factor, and $T$ is the time horizon.
At each time step $t$, each agent $n\in\{1,\ldots,N\}$ selects an action $a^{n}_{t}\in\mathcal{A}$, resulting in a joint action $\bm{a_{t}}=\{a^{1}_{t},a^{2}_{t},\ldots,a^{N}_{t}\}$. The environment transitions to a new state $s_{t+1}$ according to $\mathcal{P}(s_{t+1}|s_{t},\bm{a_{t}})$ and provides a shared reward $r(s_{t},\bm{a_{t}})$. Each agent receives an observation $z^{n}_{t}$ from $\mathcal{O}(s_{t},n)$ and maintains an observation-action history $\tau^{n}_{t}$. The goal in MARL is to learn policies $\{\pi^{n}\}_{n=1}^{N}$ that maximize the expected cumulative reward:
$J(\pi)=\mathbb{E}\left[\sum_{t=0}^{T}\gamma^{t}r(s_{t},\bm{a_{t}})\right]$ (6)
MAPPO extends PPO to multi-agent systems, optimizing the following objective function in a multi-agent context:
$L^{\text{{CLIP}}}_{\text{{MAPPO}}}(\theta)$ $=\sum_{i=1}^{n}\mathbb{E}_{t}\left[\min\left(r_{t}^{i}(\theta_{i})\hat{A}_{t},\right.\right.$ (7)
$\left.\left.\text{clip}(r_{t}^{i}(\theta_{i}),1-\epsilon,1+\epsilon)\hat{A}_{t}\right)\right]$
where
$r_{t}^{i}(\theta_{i})=\frac{\pi_{\theta_{i}}(a_{t}^{i}|o_{t}^{i})}{\pi_{\theta_{i,\text{old}}}(a_{t}^{i}|o_{t}^{i})}$ (8)
Each agent updates its policy parameters as follows:
$\theta_{i}\leftarrow\theta_{i}+\alpha\nabla_{\theta_{i}}L^{\text{CLIP}}_{i}(\theta_{i})$ (9)
MAPPO uses centralized training with decentralized execution (CTDE)[48] to handle multi-agent learning challenges. This approach maintains stability in a constantly changing environment. In MAPPO’s implementation, separate neural networks learn a policy ($\pi_{\theta}$) and a value function ($V_{\phi}(s)$) similar to the single-agent version. The value function helps reduce training variability and can take in additional global information.
This paper introduces Multi-Agent Reinforcement Learning for Single Quadruped Robot Locomotion (MASQ), which applies MARL principles to treat different parts of the robot as independent agents, trained collaboratively using shared rewards. This approach leads to faster learning in simulations and improved robustness in deployment.
IV-B MASQ parameters
Figure 4: Contact force of four legs in different gaits
State Space and Observation: The shared actor network takes a concatenated observation from four agents, each with 35 dimensions, including motor positions $q_{t}\in\mathbb{R}^{3}$, motor speeds $\dot{q_{t}}\in\mathbb{R}^{3}$, previous actions $a_{t-1}\in\mathbb{R}^{3}$ and $a_{t}\in\mathbb{R}^{3}$, a gait sequencing director $d_{t}\in\mathbb{R}^{1}$, projected gravity $g_{t}\in\mathbb{R}^{3}$, command values $v_{t}^{cmd}\in\mathbb{R}^{15}$, body speeds $v_{b}\in\mathbb{R}^{3}$, and a one-hot encoding $e_{t}\in\mathbb{R}^{1}$. The total input for the actor network is $o_{t}^{\text{actor}}\in\mathbb{R}^{140}$ (35x4). The Critic network, using global observations, has a 73-dimensional input, including motor positions $q_{t}\in\mathbb{R}^{12}$, motor speeds $\dot{q_{t}}\in\mathbb{R}^{12}$, previous actions $a_{t-1}\in\mathbb{R}^{12}$ and $a_{t}\in\mathbb{R}^{12}$, gait directors $d_{t}\in\mathbb{R}^{4}$, projected gravity $g_{t}\in\mathbb{R}^{3}$, command values $v_{t}^{cmd}\in\mathbb{R}^{15}$, and body speeds $v_{b}\in\mathbb{R}^{3}$. The output of the Critic network consists of continuous $V$ values $V_{t}\in\mathbb{R}^{4}$, which are used to calculate the advantage function.
Action Space: The output of the actor-network consists of continuous actions $a_{t}\in\mathbb{R}^{12}$, and the system then uses these to calculate the torques for the 12 motors.
Reward Function: The reward functions used in this study are designed to optimize the robot’s performance by encouraging desired behaviors and penalizing undesired ones. Key rewards include: Tracking Linear Velocity, which uses an exponential decay function to minimize velocity error; Linear Velocity Z and Angular Velocity XY, both penalizing unwanted motions to ensure stability; Torques, DOF Velocity, and DOF Acceleration, which promote energy efficiency and smoother movements; and Collision, which penalizes excessive contact forces. Additional rewards focus on action smoothness, accurate jumping, minimizing foot slip and impact velocities, and enhancing locomotion learning stability using Raibert’s heuristic and foot velocity tracking.
Figure 5: Robustness test
TABLE I: Reward settings, corresponding equations, and their values.
Reward Term Equation Value
Tracking Linear Velocity $\exp\left(-\frac{\|\mathbf{v}_{\text{cmd}}-\mathbf{v}_{b}\|^{2}}{\sigma_{t}}\right)$ 1.0
Tracking Angular Velocity $\exp\left(-\frac{\left(\mathbf{\omega}_{\text{cmd},z}-\mathbf{\omega}_{b,z}\right)^{2}}{\sigma_{\text{yaw}}}\right)$ 0.5
Linear Velocity Z $\|\mathbf{v}_{b,z}\|^{2}$ $-2\times 10^{-2}$
Angular Velocity XY $\sum_{i}\|\mathbf{\omega}_{b}\|^{2}$ $-1\times 10^{-3}$
Angular Velocity Torques $\sum_{i}\|\tau_{i}\|^{2}$ $-1\times 10^{-5}$
DOF Velocity $\sum_{i}\|\mathbf{v}_{d,i}\|^{2}$ $-1\times 10^{-4}$
DOF Acceleration $\sum_{i}\left(\frac{\mathbf{v}_{d,i,t}-\mathbf{v}_{d,i,t-1}}{\Delta t}\right)^{2}$ $-2.5\times 10^{-7}$
Collision $\sum\left(1.0\cdot(\|\mathbf{f}_{c}\|>0.1)\right)$5.0
Action Rate $\sum\|\mathbf{a}_{t}-\mathbf{a}_{t-1}\|^{2}$ $-1\times 10^{-2}$
Jump $-(h_{b}-h_{j})^{2}$ 10.0
Feet Slip $\sum(c_{f}\cdot\|\mathbf{v}_{f}\|^{2})$ $-4\times 10^{-2}$
Action Smoothness 1 $\sum\left(\mathbf{a}_{t}-\mathbf{a}_{t-1}\right)^{2}$0.1
Action Smoothness 2 $\sum\left(\mathbf{a}_{t}-2\mathbf{a}_{t-1}+\mathbf{a}_{t-2}\right)^{2}$0.1
Feet Impact Velocity $\sum(c_{s}\cdot\|\mathbf{v}_{f,p}\|^{2})$0.0
Raibert Heuristic $\sum\|e_{r}\|^{2}$10.0
Tracking Contacts Shaped Velocity $\sum\left(c_{d}\cdot\left(1-\exp\left(-\frac{\|\mathbf{v}_{f}\|^{2}}{\sigma_{gv}}\right)\right)\right)$ 4.0
IV-C Multi-agent actor and global critic networks
In the simulated environment of Isaac Gym, the robot receives observations and rewards to facilitate its learning. The learning process involves dividing the observations into shared and private features, and both the actor and critic networks use the rewards to train the policy.
The actor-network consists of a multi-layer perception (MLP) base and an activation layer that produces actions and their associated log probabilities. Similarly, the critic network uses an MLP base, ending in an output layer that predicts value functions. The actor-network selects actions based on observations, while the critic network assesses the value of these actions to guide the learning process. The actions produced for the agents are processed by an actuator network to simulate real-world conditions, enhancing deployment effectiveness in natural environments. After training, the trained actor-network is deployed onto the robotic dog to perform actions directly. Fig.3 illustrates the entire learning process.
In the context of a quadruped robot, we consider each leg as an individual agent. All four agents share a standard actor-network. Using a shared-parameter network instead of four separate actor networks helps reduce computational load and better fits the nature of the quadruped robot. Unlike typical multi-agent environments, such as Starcraft, where each soldier is an independent agent, the quadruped robot is a single entity with four legs symmetrically positioned around the body’s center. Therefore, a shared-parameter actor-network is more suitable for this scenario.
Because all four legs share the same actor network, we express the policy for each leg as follows:
$\pi_{\theta_{i}}(a_{i,t}\mid s_{i,t})=\pi_{\theta}(a_{i,t}\mid s_{i,t})$ (10)
where i = 1, 2, 3, 4 corresponds to the four legs.
The quadruped robot has four legs, each with three motors: Hip, Thigh, and Calf joint motors. Each motor’s position, velocity, and torque are observable, so we use these details as the independent observations for each agent. Additionally, to enhance the coordination among the agents, we augment each independent observation with shared observations, including speed and gravity projections calculated from IMU data, temporal director, and agent ID. The temporal director guides the gait sequence of each leg under different movement postures, while the agent ID is necessary for the shared-parameter network. This setup ensures the independence of each agent while improving their cooperative capabilities. The temporal director helps to synchronize the movements of different legs, ensuring smooth gait patterns. It can be defined as:
$T_{i}(t)=\sin\left(2\pi\times F_{i}(t)\right)$ (11)
$F_{i}(t)=G(t)+\phi_{i}+\Delta_{i}$ (12)
where:
  • $F_{i}(t)$ represents the temporal index of the $i$-th leg at time $t$.
  • $G(t)$ is the current phase of the gait cycle, increasing over time and wrapping within $[0,1)$.
  • $\phi_{i}$ is the phase offset associated with the gait for the $i$-th leg.
  • $\Delta_{i}$ is an additional time offset for the $i$-th leg, which is related to the gait.
The Global Critic method uses a centralized value function approach to consider global information, which fits into the CTDE. It relies on a worldwide value function to coordinate individual agents, such as single-agent algorithms like PPO. The Critic network takes in global observations, which consist of the separate observations of the four agents and shared global information.
During the training in the simulation environment, we used a multi-gait curriculum learning[49]. This curriculum comprises four gaits: pace, trot, bound, and pronk. In the simulation, the quadruped robot learns these four gaits simultaneously, and the progress is updated based on evaluating whether the gaits’ rewards meet specific thresholds. This method allows the robot to learn different gaits effectively and is also helpful in testing the generalization capabilities of our proposed approach across various tasks in experimental settings.
IV-D Sim-to-real
To bridge the gap between simulation and real-world performance, we utilized domain randomization and an Actuator Network. Domain randomization involves randomizing various parameters to train a robust policy under different conditions[27, 50, 51]. These parameters include robot body mass, motor strength, joint position calibration, ground friction, restitution, orientation, and magnitude of gravity. We also independently randomize friction coefficients, including ground friction and tile height. Gravity is randomized every 8 seconds with a gravity impulse duration of 0.99 seconds. The system randomizes the time steps every 6 seconds and performs the overall randomization every 4 seconds. These measures enhance the model’s robustness and adaptability. The training process for the Actuator Network captures the non-ideal relationship between PD error and the torque realized [52], thereby improving the model’s performance in real-world applications.
V Experiments
We conducted experiments using the Unitree Go2 quadrupled robot with various experiments. Section V-A introduces the experimental setup. Section V-C1 shows the experimental results of the simulation. Section V-C shows the real-world experiments and comparisons.
V-A Experiments setup
In the simulation environment, we have designed two types of terrain: flat and uneven. Uneven terrain is a composite of pyramid-sloping terrain and random uniform terrain. Fifteen commands, sampled within a specified range, guide the gait-learning process. The robot learns to adapt and develop various skills by following these commands.
V-B Experiments in simulation
Figure 6: Flat Terrain
Figure 7: Uneven Terrain
Figure 8: Reward over Time (Flat Terrain)
Figure 9: Reward over Time (Uneven Terrain)
We compared the proposed learning algorithm with the PPO student-teacher baseline in simulation environments. We set up two types of simulation environments: flat and uneven terrain. We trained the quadruped robot using the curriculum learning method in both environments. The reward curves show that the proposed method achieves faster convergence and superior final performance compared to the PPO baseline. We present the experimental results as the mean of the rewards over five tests for a fair comparison.
V-C Real-word experiments
This paper deploys MASQ on a quadruped robot and tests it on various terrains, including flat ground, grass, and sand, where it performs excellently. In addition, we conducted a series of robustness tests for heavy impacts and side kicks. The robot can quickly return to its normal state after being disturbed.
V-C1 Gaits test
Fig. 4 illustrates the periodic relationship of quadruped force feedback for each of the four legs under the training of four different skills. They reflect the impact of incorporating the temporal director in our observations on the learning and switching of other skills.
V-C2 Robustness test
To validate the robustness of the gaits trained with our method, we conduct external disturbance tests on the robot. We test the robot in a bounding gait with a cycle period of 20ms. During robustness tests, the robot performs continuous jumps in this bounding gait. The impacts of human steps on the robot are applied to introduce disturbances during its jump cycle. Under normal conditions, the robot landed simultaneously on all four legs. We monitor this process by recording the force sensor values in the robot’s feet, thereby documenting the transition from a normal state through the induced disturbance and back to the normal state. As shown in Fig. 5, the robot returns to its normal state after experiencing a disturbance within just one gait cycle.
VI Conclusion and Future Work
This paper proposes a multi-agent-based motion control system for quadruped robots, leveraging a shared-parameter actor network and a centralized critic network within the CTDE framework. Our approach, implemented in Isaac Gym, demonstrates significant improvements in training speed and robustness through curriculum learning and domain randomization, leading to efficient limb cooperation and smooth sim-to-real transitions. Experiments validate the effectiveness of the CTDE method in enhancing cooperation and efficiency in motion control for symmetric robots.
Future work will extend this approach to other symmetric robots and explore advanced domain adaptation techniques to further minimize the sim-to-real gap, enhancing performance in dynamic environments.